I do not fully sure get why (-4)squared is 16, but -4 squared without perenthisis is -16. In other words, (-A)squared =A squared, but -A squared = -A squared. Please clear this up! Also, I do not see how to show exponents on the computer. Thanks, Drew Lurie
Answer by mathrebel(25)

power is ^
2^2=4
3^2=9
This is a theoretical question.
(-4)^2 means that u square whatever is inside the parenthesis, so the answer is (-4)*(-4)=16
-4^2 means that u only square the number, so the answer is:
-(4)*(4)=-16.